OpenStudy (anonymous):

Question...How would I graph the domain of the following function on a number line?

OpenStudy (anonymous):

\[g(x)=\ln(5+4x-x^2)\]\[g(x)=\ln[(-x+5)(x+1)]\]the function is not defined for for x=5 and x=-1; there will be a vertical asymptote there.\[g'(x)=\frac{2(x-2)}{(x-5)(x+1)}\]the function will be horizontal at x=2. the function is not defined for values x>=5 and x<=-1 the graph will end up looking like this